# Teaching LLMs According to Their Aptitude: Adaptive Reasoning for Mathematical Problem Solving

### v2

@@ Abstract
Existing approaches to mathematical reasoning with large language models (LLMs) rely on Chain-of-Thought (CoT) for generalizability or Tool-Integrated Reasoning (TIR) for precise computation. While efforts have been made to combine these methods, they primarily rely on post-selection or predefined strategies, leaving an open question: whether LLMs can autonomously adapt their reasoning strategies based on their inherent capabilities. In this work, we propose TATA (Teaching LLMs According to Their Aptitude), an adaptive framework that enables LLMs to personalize their reasoning strategy spontaneously, aligning it with their intrinsic aptitude. TATA incorporates base-LLM-aware data selection during supervised fine-tuning (SFT) to tailor training data to the model’s unique abilities. This approach equips LLMs to autonomously determine and apply the appropriate reasoning strategy at test time. We evaluate TATA through extensive experiments on six mathematical reasoning benchmarks, using both general-purpose and math-specialized LLMs. Empirical results demonstrate that TATA effectively combines the complementary strengths of CoT and TIR, achieving superior or comparable performance with improved inference efficiency compared to TIR alone. Further analysis underscores the critical role of aptitude-aware data selection in enabling LLMs to make effective and adaptive reasoning decisions and align reasoning strategies with model capabilities.
@@
Xin Xu1,2, Yan Xu1,2, Tianhao Chen1, Yuchen Yan3, Chengwu Liu2,4, Zaoyu Chen2,5, Yufei Wang2, Yichun Yin2, Yasheng Wang2, Lifeng Shang2, Qun Liu2, 1The Hong Kong University of Science and Technology, 2Huawei Noah’s Ark Lab, 3Zhejiang University, 4Peking University, 5The Hong Kong Polytechnic University Correspondence: {xxuca, yxucb}@connect.ust.hk
**footnotetext: Work done during the internship at Huawei Noah’s Ark Lab. Codes and data are available at https://github.com/XinXU-USTC/TATA.
@@ 1 Introduction
Mathematical reasoning is a cornerstone of human intelligence, and enabling machines to solve mathematical problems is crucial for many applications (Ahn et al., 2024; He et al., 2024). Recent years have seen rapid progress in leveraging LLMs for mathematical reasoning, with advancements in continuing pretraining (Lewkowycz et al., 2022; Azerbayev et al., 2023; Shao et al., 2024), supervised fine-tuning (SFT) (Yu et al., 2023; Tong et al., 2024; Yan et al., 2024; Toshniwal et al., 2024), reinforcement learning (RL) (Shao et al., 2024; DeepSeek-AI et al., 2025), prompting (Wei et al., 2022; Zheng et al., 2023; Wang et al., 2023b), and evaluation (Tang et al., 2024; Xu et al., 2025b; Gao et al., 2024; Xu et al., 2025a). Among these, SFT has attracted significant attention, even in recent test-time scaling efforts (Muennighoff et al., 2025).
@@
LLM math reasoning typically follows two paradigms: Chain-of-Thought (CoT) (Wei et al., 2022), which uses natural language intermediate steps for flexibility and interpretability but lacks computational guarantees, and Tool-Integrated Reasoning (TIR) (Chen et al., 2022; Gao et al., 2023), which employs executable code for precise computations at the cost of higher resource demands due to interactive execution. In contrast, CoT requires only a single forward pass, making it more computationally efficient. Previous work often trains LLMs exclusively on CoT/TIR data (Tong et al., 2024; Shao et al., 2024; Yan et al., 2024; Gou et al., 2023; Wang et al., 2023a; Lu et al., 2024), or relies on external signals or predefined prompts to select between the two (Zhao et al., 2023; Yue et al., 2023; Yang et al., 2024b) (see Figure 1). Recent advancements in slow-thinking models, such as DeepSeek R1 (DeepSeek-AI et al., 2025), have demonstrated the potential of long CoT reasoning. Then a promising direction lies in enabling LLMs to adaptively incorporate tools during slow-thinking processes, combining the strengths of both CoT and TIR (Figure 1 (d)). A critical first step toward this goal is to teach base LLMs to adaptively select between CoT and TIR according to their inherent aptitude through SFT. By doing so, we can better unlock the model’s potential to use tools effectively while maintaining its generalization ability across diverse problem-solving scenarios (Figure 1 (d))
@@
In this work, we propose Teaching LLMs According to Their Aptitude (TATA), an adaptive framework that enables LLMs to spontaneously select between CoT and TIR for math problem solving by using SFT data selection that aligns with LLM’s inherent aptitude. Unlike previous methods, TATA leverages base-LLM-aware data selection during SFT, ensuring that the training process not only suits the model’s potential but also mitigates the risk of catastrophic forgetting (Chu et al., 2025). By adaptively selecting between CoT and TIR based on the base model’s aptitude, TATA preserves and enhances the generalizability of the model, particularly for out-of-domain tasks.
@@
Concretely, we begin with a dataset $\mathcal{D}$, which consists of $N$ triplets, each containing a query, a CoT solution, and a TIR solution. We then construct an anchor set, $\mathcal{D}_{\text{anchor}}$, to evaluate the model’s performance. For each training query in $\mathcal{D}$, we assess the LLM’s accuracy on $\mathcal{D}_{\text{anchor}}$ by providing either the CoT or TIR solution of the query as a one-shot example. Based on the model’s performance in each setting, we select the most effective reasoning paradigm for training queries and use it to construct the SFT data, $\mathcal{D}_{\text{SFT}}$.
To assess the effectiveness of TATA, we conduct extensive evaluations across six mathematical reasoning benchmarks, utilizing both general-purpose LLMs (e.g. Llama-3-8B (AI@Meta, 2024)) and math-specialized LLMs (e.g. Qwen2.5-Math-7B (Yang et al., 2024b)) as base models. Experimental results demonstrate that TATA successfully personalizes reasoning strategies according to each LLM’s aptitude, leading to better performance across different models and benchmarks.
@@ 3.1 Problem Setting
Many advanced capabilities can be distilled through SFT (Muennighoff et al., 2025; Zhou et al., 2023b; DeepSeek-AI et al., 2025). Building on this, we aim to teach LLMs to spontaneously select between CoT and TIR using SFT by carefully selecting reasoning patterns tailored to different training queries. In this section, we formally formulate our research question as an SFT data selection problem.
@@
Our objective is to construct an SFT dataset from the candidate dataset $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$ by incorporating suitable reasoning patterns for different training queries. Specifically, for each problem $x_{i}$ in $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$, we need to decide whether to include its CoT solutions or TIR solutions in the SFT dataset. Formally, this involves determining whether $\{(x_{i},y_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}}$ or $\{(x_{i},z_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}}$.***We also consider scenarios where both CoT and TIR solutions for a query are included in the SFT dataset. For example, CoT-only SFT (Xu et al., 2024c) constructs the dataset such that $\{(x_{i},y_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}},\forall i$. In contrast, TIR-only SFT (Gou et al., 2023) selects $\{(x_{i},z_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}},\forall i$. Our work aims to go beyond static selection approaches by dynamically tailoring the reasoning paradigm to suit the specific requirements of each training query, while also accounting for the base LLM’s aptitude.
@@ 3.2 TATA Overview
Following the success of distillation through SFT (Muennighoff et al., 2025; DeepSeek-AI et al., 2025), we aim to enable LLMs to adaptively select appropriate reasoning strategies based on their own aptitude, guided by the control of SFT data selection (see Section 3.1). Intuitively, if an LLM demonstrates improved performance on certain queries when trained with CoT solutions instead of TIR solutions, it suggests its inclination toward CoT reasoning in those cases. This preference can be extrapolated to new cases, where the model is expected to favor CoT for similar problems during testing. The same principle applies to TIR-based reasoning. Inspired by IIT theory (see Section 2.3), LLMs can be indirectly “fine-tuned” with CoT or TIR examples through one-shot learning, thereby replacing the need for actual SFT.
@@
As depicted in Figure 2, our proposed framework, TATA, comprises four main steps: data construction, anchor construction, contribution quantification, and data selection. In the data construction stage, we adapt an original training set, $\mathcal{D}_{\text{orig}}$, containing CoT solutions, to form the candidate set $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$. This candidate set includes triplets of queries, a CoT solution, and corresponding TIR solution. Next, during the anchor construction stage, a representative anchor set of size $A$ is generated from the original training set by clustering. In the contribution quantification stage, we compute two scores, ${S_{\text{CoT}}^{k}}$ and $S_{\text{TIR}}^{k}$, for each query $q_{k}$ in the candidate set $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$. These scores indicate the impact of CoT and TIR solutions on the performance of LLMs using IIT (see Setion 2.3). Finally, the data selection step formulates a decision based on ${S_{\text{CoT}}^{k}}$ and $S_{\text{TIR}}^{k}$, determining whether to include CoT or TIR solutions for queries in $\mathcal{D}$.
@@ 3.3 TATA Details
To evaluate the impact of specific CoT or TIR solutions on the performance of LLMs, we construct an anchor set, denoted by $\mathcal{D}_{\text{anchor}}=\{(q_{i},a_{i})\}_{i=1}^{A}$, where $A$ is the size of the anchor set, $q_{i},a_{i}$ is the $i$-th question and corresponding ground-truth answer in $\mathcal{D}_{\text{anchor}}$. We expect $\mathcal{D}_{\text{anchor}}$ to be diverse, ensuring that accuracy on this set fairly reflects the LLMs’ overall performance. To achieve this, we first encode all queries from $\mathcal{D}_{\text{orig}}$ into vector representations using an embedding model (e.g., text-embedding-ada-002) and then cluster them into $A$ distinct groups. The center of each cluster is selected to $\mathcal{D}_{\text{anchor}}$. This approach takes the semantic diversity of questions into account, making $\mathcal{D}_{\text{anchor}}$ a reliable indicator of LLMs’ performance.
@@
LLaMA-3-8B CoT 84.7 46.5 65.6 91.6 81.6 30.2 13.3 54.2 58.0
TIR 81.7 56.2 69.0 87.8 77.8 30.5 21.9 54.5 59.3
@@
Qwen2.5Math-1.5B CoT 83.1 56.6 69.8 92.9 84.1 44.3 19.3 60.2 63.4
TIR 78.8 64.8 71.8 92.3 83.3 42.4 27.4 61.4 64.8
@@
Qwen2.5Math-7B CoT 91.0 61.5 76.2 94.8 87.9 45.7 23.9 63.1 67.5
TIR 88.9 73.6 81.2 95.4 89.4 47.1 35.3 66.8 71.6
@@
Table 1: The accuracies (%) of our TATA framework, comparing with CoT and TIR methods. The best accuracies within each group are shown in bold. “ID AVG”, “OOD AVG”, and “AVG” denote the averages of these metrics across in-domain, out-of-domain, and all six benchmarks. “CoT”, “TIR”, and “TATA” indicate fine-tuning exclusively on CoT data, TIR data, and using our TATA framework, respectively, with the corresponding base LLM.

### v3

@@ Abstract
Existing supervised fine-tuning (SFT) approaches to enhance the mathematical reasoning of large language models (LLMs) rely either on Chain-of-Thought (CoT) for generalizability or Tool-Integrated Reasoning (TIR) for precise computation. While efforts have been made to combine these methods, they primarily rely on post-selection or predefined strategies, leaving an open question: Could we endow LLMs with the ability to adaptively determine whether to use CoT or TIR based on the math problems at hand? In this work, we propose TATA (Teaching LLMs According to Their Aptitude), an adaptive framework that enables LLMs to personalize their reasoning strategy for different problems spontaneously, aligning it with their intrinsic aptitude. TATA incorporates base-LLM-aware data selection during SFT to tailor training data to the model’s unique abilities, which equips LLMs to autonomously determine and apply the effective reasoning strategy at test time. Empirical results demonstrate that TATA effectively combines the complementary strengths of CoT and TIR, achieving superior or comparable performance with improved inference efficiency compared to existing methods. Further analysis underscores the critical role of aptitude-aware data selection in enabling LLMs to make effective and adaptive reasoning decisions and align reasoning strategies with model capabilities.
@@
Xin Xu1, Yan Xu2, Tianhao Chen1, Yuchen Yan3, Chengwu Liu4, Zaoyu Chen5, Yufei Wang2, Yichun Yin2, Yasheng Wang2, Lifeng Shang2, Qun Liu2 1The Hong Kong University of Science and Technology, 2Huawei Noah’s Ark Lab, 3Zhejiang University, 4Peking University, 5The Hong Kong Polytechnic University xxuca@connect.ust.hk, xuyan110@huawei.com
footnotetext: * Work done during the internship at Huawei Noah’s Ark Lab. Corresponding author. Codes and data are available at https://github.com/XinXU-USTC/TATA.
@@ 1 Introduction
Previous SFT methods for mathematical reasoning (Tong et al., 2024; Shao et al., 2024; Yan et al., 2024; Gou et al., 2023; Wang et al., 2023; Lu et al., 2024) predominantly adopt one of the following two distinct reasoning paradigms: Chain-of-Thought (CoT) reasoning (Wei et al., 2022) or Tool-Integrated Reasoning (TIR) (Chen et al., 2022; Gao et al., 2023). CoT employs natural language (NL) to articulate intermediate reasoning steps, whereas TIR integrates NL with Python code blocks in an interleaved manner (see Section 3.2). While CoT offers computational efficiency, it may compromise the numerical accuracy of complex calculations. In contrast, TIR’s structured execution of code ensures precise computation but incurs significant computational overhead. Notably, recent studies (Zhao et al., 2023; Yang et al., 2024b) have empirically demonstrated that CoT and TIR exhibit complementary strengths: CoT demonstrates superior performance on problems demanding sophisticated logical deduction with minimal numerical computation, whereas TIR excels in scenarios requiring intensive numerical calculations with relatively simpler logical flow.
This complementary nature suggests potential benefits to integrate these two reasoning patterns. Zhao et al. (2023) proposes an auxiliary LLM-based selector to dynamically choose between paradigms via prompt-based routing (Figure 1 (a)). MAmmoTH (Yue et al., 2023) switches to CoT reasoning if TIR encounters execution errors or timeouts (Figure 1 (b)). Yang et al. (2024b) employs different inference prompts to elicit respective reasoning capabilities (Figure 1 (c)). Despite these advancements, existing approaches predominantly rely on either explicit external selectors (as in Zhao et al. (2023)) or predefined heuristics (as in MAmmoTH and Qwen-2.5-Math) rather than endowing LLMs with the intrinsic capability to autonomously recognize and apply appropriate reasoning strategies. However, the potential for LLMs themselves to dynamically adapt reasoning paradigms (CoT or TIR) remains underexplored.
@@
To bridge this gap, we propose Teaching LLMs According to Their Aptitude (TATA), an adaptive framework that enables LLMs to spontaneously select between CoT and TIR for math problem solving. Instead of adopting a fixed strategy for all training queries, TATA adaptively tailors the training data selection process by considering both the query characteristics and the base LLM’s aptitude. This ensures that the resulting model is equipped to select a suitable reasoning strategy (CoT or TIR) for different queries at test time, facilitating aptitude-driven reasoning. As a result, TATA preserves and enhances the generalizability of the model, particularly for out-of-domain tasks.
@@
Concretely, we begin with a dataset $\mathcal{D}$, which consists of $N$ triplets, each containing a query, a CoT solution, and a TIR solution. We then construct an anchor set, $\mathcal{D}_{\text{anchor}}$, to evaluate the model’s performance. For each training query in $\mathcal{D}$, we assess the LLM’s accuracy on $\mathcal{D}_{\text{anchor}}$ by providing either the CoT or TIR solution of the query as a one-shot example. Based on the model’s performance on the $\mathcal{D}_{\text{anchor}}$ in each setting, we select the most effective reasoning paradigm for training queries and use it to construct the SFT data, $\mathcal{D}_{\text{SFT}}$. We endow the base LLMs with the ability to adaptively switch between CoT and TIR by training of personalized training set $\mathcal{D}_{\text{SFT}}$. To assess TATA’s effectiveness, we conduct extensive evaluations across six math reasoning benchmarks, utilizing both general-purpose LLMs (e.g. Llama-3-8B (AI@Meta, 2024)) and math-specialized LLMs (e.g. Qwen2.5-Math-7B (Yang et al., 2024b)) as base models. Experimental results show that TATA successfully leads to better performance across different models and benchmarks.
@@
2 Related Work
Math Reasoning with CoT and TIR
CoT and TIR are two widely recognized approaches for reasoning with LLMs. CoT offers interpretability and generalizability, while TIR can provide precise calculation results. Previous work on mathematical SFT has primarily focused on either CoT (Yu et al., 2023; Tong et al., 2024; Shao et al., 2024; Yan et al., 2024) or TIR (Yue et al., 2023; Gou et al., 2023; Wang et al., 2023; Yin et al., 2024), with a few efforts to integrate both (Yue et al., 2023; Beeching et al., 2024; Yang et al., 2024b). For instance, MAmmoTH (Yue et al., 2023) mainly adopts TIR but switches to CoT when code execution fails due to errors or timeouts. However, it relies on separate prompts and manual inference controls to switch between them. Recent work has explored automatic selection between CoT and TIR (Zhao et al., 2023; Yue et al., 2024; Yu et al., 2024), such as using an auxiliary LLM to determine CoT/TIR (Zhao et al., 2023). However, these methods rely on external planners to select CoT/TIR, not by LLMs themselves. In contrast, our work seeks to enable LLMs to spontaneously select the appropriate reasoning strategy without relying on external planners or manual interventions.
Data Selection
Test-Time Scaling
@@ 4.1 Problem Setting
In this section, we formally formulate our problem setting, including our data structure and objective.
@@
Our objective is to construct an SFT dataset from the candidate dataset $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$ by incorporating suitable reasoning patterns for different training queries. Specifically, for each problem $x_{i}$ in $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$, we need to decide whether to include its CoT solutions or TIR solutions in the SFT dataset. Formally, this involves determining whether $\{(x_{i},y_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}}$ or $\{(x_{i},z_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}}$. [^1] We also consider scenarios where both CoT and TIR solutions for a query are included in the SFT dataset. For example, CoT-only SFT (Xu et al., 2024c) constructs the dataset such that $\{(x_{i},y_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}},\forall i$. In contrast, TIR-only SFT (Gou et al., 2023) selects $\{(x_{i},z_{i}^{j})\}_{j=1}^{M_{i}}\subseteq D_{\text{SFT}},\forall i$. Unlike these static selection approaches, TATA aims to dynamically tailor the most suitable reasoning paradigm for different training queries and base LLMs.
@@ 4.2 TATA Overview
Intuitively, if an LLM demonstrates improved performance on certain queries when fine-tuned with CoT solutions instead of TIR solutions, it suggests its inclination toward CoT reasoning in those cases. This preference can be extrapolated to new cases, where the model is expected to favor CoT for similar problems during testing. The same principle applies to TIR-based reasoning. Inspired by IIT theory (see Section 3.3), LLMs can be indirectly “fine-tuned” with CoT or TIR examples through one-shot learning, thereby replacing the need for actual SFT.
@@
As depicted in Figure 2, our proposed framework, TATA, comprises four main steps: data construction, anchor construction, contribution quantification, and data selection. In the data construction stage, we adapt an original training set, $\mathcal{D}_{\text{orig}}$, containing CoT solutions, to form the candidate set $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$. This candidate set includes triplets of queries, a CoT solution, and corresponding TIR solution. Next, during the anchor construction stage, a representative anchor set of size $A$ is generated from the original training set by clustering. In the contribution quantification stage, we compute two scores, ${S_{\text{CoT}}^{k}}$ and $S_{\text{TIR}}^{k}$, for each query $q_{k}$ in the candidate set $\mathcal{D}={\{(x_{i},y_{i}^{j},z_{i}^{j})\}_{i=1}^{N}}_{j=1}^{M_{i}}$. These scores indicate the impact of CoT and TIR solutions on the performance of LLMs using IIT (see Section 3.3). The data selection step formulates a decision based on ${S_{\text{CoT}}^{k}}$ and $S_{\text{TIR}}^{k}$, determining whether to include CoT or TIR solutions for queries in $\mathcal{D}$. Finally, SFT is performed on this curated training set.
@@ 4.3 TATA Details
To evaluate the impact of specific CoT or TIR solutions on the performance of LLMs, we construct an anchor set, denoted by $\mathcal{D}_{\text{anchor}}=\{(q_{i},a_{i})\}_{i=1}^{A}$, where $A$ is the size of the anchor set, $q_{i},a_{i}$ is the $i$-th question and corresponding ground-truth answer in $\mathcal{D}_{\text{anchor}}$. We expect $\mathcal{D}_{\text{anchor}}$ to be diverse, ensuring that accuracy on this set fairly reflects the LLMs’ overall performance. To achieve this, we first encode all queries from $\mathcal{D}_{\text{orig}}$ into vector representations using an embedding model (e.g., text-embedding-ada-002) and then cluster them into $A$ distinct groups. The center of each cluster is selected to $\mathcal{D}_{\text{anchor}}$. This approach takes the semantic diversity of questions into account, making $\mathcal{D}_{\text{anchor}}$ a reliable indicator of LLMs’ performance. To put it simple, one can treat this $\mathcal{D}_{\text{anchor}}$ as a validation set.
@@
Qwen2.5-14B hybrid 91.4 71.7 81.5 93.8 84.5 45.8 35.3 64.8 70.4
ensemble 90.1 66.9 78.5 92.2 82.8 46.1 32.3 63.3 68.4
GPT-Select 90.7 61.5 76.1 86.2 79.1 44.1 23.0 58.1 64.1
TATA 92.1 71.7 81.9 96.5 88.4 46.4 35.3 66.7 71.7
LLaMA-3-8B hybrid 82.0 56.1 69.1 88.0 78.0 30.8 21.3 54.5 59.4
ensemble 84.0 46.9 65.4 88.6 79.3 29.6 15.3 53.2 57.3
GPT-Select 83.2 47.2 65.2 85.3 77.5 30.6 13.9 51.8 56.3
@@
Qwen2.5Math-1.5B hybrid 82.6 66.3 74.4 92.7 83.6 43.1 26.2 61.4 65.7
ensemble 81.5 64.7 73.1 91.8 83.9 44.1 27.4 61.8 65.6
GPT-Select 79.4 56.9 68.1 92.7 83.7 41.8 20.6 59.7 62.5
@@
Qwen2.5Math-7B hybrid 89.2 73.4 81.3 95.4 89.5 47.1 34.4 66.6 71.5
ensemble 89.1 67.7 78.4 93.4 84.5 46.7 30.8 63.9 68.8
GPT-Select 89.8 63.0 76.4 89.4 85.1 44.4 24.6 60.7 65.9
@@
Table 1: The accuracies (%) of our TATA framework, comparing with various baselines. The best accuracies within each group are shown in bold. “ID AVG”, “OOD AVG”, and “AVG” denote the averages of these metrics across in-domain, out-of-domain, and all six benchmarks.
@@
Model Method In-Domain Out-of-Domain AVG
GSM8K MATH ID AVG MAWPS SVAMP College Olympiad OOD AVG
Qwen2.5-0.5B hybrid 49.3 37.7 43.5 84.5 55.0 27.5 7.9 43.7 43.6
ensemble 47.1 34.8 41.0 83.4 53.8 25.6 7.7 42.6 42.1
GPT-Select 45.6 31.6 38.6 80.4 52.6 24.4 7.1 41.1 40.3
CoT+TIR 51.5 33.5 42.5 85.8 58.6 25.7 7.9 44.4 43.8
TATA - random 100 50.6 34.6 42.6 85.7 57.6 26.2 6.2 43.9 43.5
TATA - A 200 52.6 36.8 44.7 85.1 59.6 27.4 8.4 45.1 45.0
TATA 52.8 36.6 44.7 85.9 59.4 26.9 8.6 45.2 45.0
Table 3: Ablation of Contribution Quantification.